Partner Bonus

Lilly has a capital balance of $113,000 after adjusting assets to fair market value. Marler contributes $60,000 to receive a 40% interest in a new partnership with Lilly.

Determine the amount and recipient of the partner bonus.

Amount of bonus $
Recipient of bonus
Answer #1

ANS :-

Total Capital of new partnership business = 113000 + 60000 = $ 173000

Interest of Marler = 173000 * 40 % = $ 69200

Bonus amount = 69200 - 60000 = $ 9200

The Bonus of $ 9200 will be paid Lilly.

Conclusion:-

Amount of Bonus = $ 9200

Recipient of Bonus = Lilly.